OpenStudy (anonymous):

Let\[L(x)=\int_{1}^{x}\frac{1}{t}dt.\]How can I show that\[L(ab)=L(a)+(b)?\]

OpenStudy (kinggeorge):

The key to this, would be to notice that the integral of 1/t is the natural log, and that property holds with natural log.

OpenStudy (anonymous):

I only want to use integral properties. Otherwise, this problem would be trivial.

OpenStudy (anonymous):

perhaps you can start by showing that \[\int_1^{2x}\frac{dt}{t}=\int_1^2\frac{dt}{t}+\int_1^x\frac{dt}{t}\]

OpenStudy (anonymous):

take the derivatives, show that they are the same, therefore the functions must differ by a constant, find the constant

OpenStudy (anonymous):

first solve for the integral in the LHD so that will give you \[L(x)= \int\limits_{1}^{x} 1/t dt =\ln\left| x \right|-\ln\left| 1 \right|=\ln\left| x \right|\] so since: \[ L(x)=\ln \left| x \right|\] \[ L(ab)=\ln\left| ab \right|=\ln\left| a\right|\left| b \right|=\ln\left| a \right|+\ln\left| b \right|=L(a)+L(b)\] therefore, L(ab)=L(a)+L(b) notice i used a property of logs

OpenStudy (anonymous):

in my example \[\frac{d}{dx}\int_1^{2x}\frac{dt}{t}=\frac{1}{x}\] \[\frac{d}{dx}\int_1^x\frac{dt}{t}=\frac{1}{x}\] therefore \[\int_1^{2x}\frac{dt}{t}=\int_1^{x}\frac{dt}{t}+C\]

OpenStudy (anonymous):

and you can find C easily

OpenStudy (experimentx):

\[ \int_{1}^{ab} 1/x dx = \int_{1}^{a}1/x dx + \int_{a}^{ab}1/x dx \] let x = at, when x=a, t=1, and when x =ab, t=b, dx = a dt \[ \int_{a}^{ab} 1/x dx = \int_{1}^{b} adt/at = \int_{1}^{b} dt/t \] We have, \[ \int_{1}^{a}1/x dx + \int_{1}^{b}1/t dt = \int_{1}^{a}1/x dx + \int_{1}^{b}1/x dx \]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

or if you like @pre-algebra : \[ L(ab)=\int\limits_{1}^{ab} 1/t dt=\ln \left| ab \right|-\ln\left| 1 \right|=\ln\left| ab \right|=\ln\left| a \right|\left| b \right|=\ln\left| a \right|+\ln\left| b \right|\] \[=\ln\left| a \right|-0+\ln\left| b \right|-0=\ln\left| a \right|-\ln\left| 1 \right|+\ln\left| b \right|-\ln\left| 1 \right|=\int\limits_{1}^{a}1/tdt+\int\limits_{1}^{b}1/tdt\] \[=L(a)+L(b)\] so L(ab)=L(a)+L(b)

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

@satellite73, in this example, is it possible to find \(C\), or is that just a generalization?

OpenStudy (anonymous):

just an example.

OpenStudy (anonymous):

\[f(x)=\int_1^{x}\frac{dt}{t}\] \[f(ax)=\int_1^{ax}\frac{dt}{t}\] \[f'(ax)=f'(x)\] is an easy check. therefore \[f(ax)=f(x)+C\] now to find C, we know that \[f(1)=0\] so \[f(a)=f(a\times 1)=f(1)+C=C\] hence \(C=f(a)\) and thefore \[f(ax)=f(a)+f(x)\]

OpenStudy (anonymous):

Just for the record, the way I ended up doing this was by substitution. For this example, we have that\[L(ab)=\int_{1}^{a}\frac{dt}{t}+\int_{a}^{ab}\frac{dt}{t}.\]If we let \(t=au\), then \(dt=adu\), and\[\int_{a}^{ab}\frac{dt}{t}=\int_{1}^{b}\frac{adu}{au}=\int_{1}^{b}\frac{du}{u}.\]Hence,\[L(ab)=\int_{1}^{a}\frac{dt}{t}+\int_{1}^{b}\frac{du}{u}=L(a)+L(b).\]